# Robust Neural Information Retrieval: An Adversarial and Out-of-distribution Perspective

### v1

@@ 2.2. Taxonomy of Robustness in IR
In IR, robustness threats exist in a wide range of aspects, including adversarial robustness (Wu et al., 2022b), OOD robustness (Thakur et al., 2021) and performance variance (Wu et al., 2022b). The framework of robustness in IR is shown in Figure 4.
2.2.1. Performance variance
Typically, the performance of IR models is first represented by their overall performance on IID data. Recently, it has been recognized that performance stability across IID queries may be compromised when we try to improve the average retrieval effectiveness across all queries (Zhang et al., 2013). Therefore, a robust neural IR model should not only have good retrieval performance on the overall testing queries, but also ensure that the performance on individual queries is not too bad. We give a formal definition of performance variance in IR based on Definition 2.1.
Definition 2.0 (Performance variance of information retrieval).
Given an IR model $f_{\mathcal{D}_{\mathrm{train}}}$ trained on training dataset $\mathcal{D}_{\mathrm{train}}$ with a corresponding IID testing dataset $\mathcal{D}_{\mathrm{test}}$, and an acceptable error threshold $\delta$, for the top-$K$ ranking result, if
@@ 2.2.1. Performance variance
where $\mathrm{Var}(\cdot)$ is the variance of the ranking performance of the IR model $f_{\mathcal{D}_{\mathrm{train}}}$ on $\mathcal{D}_{\mathrm{test}}$, then the model $f$ is considered $\delta$-robust in terms of performance variance for metric $M$.
2.2.2. Out-of-distribution robustness
IR models need to cope with a constant stream of unseen data (Wu et al., 2022b). The key behind this challenge is how to adapt the model to new data out of the familiar distribution. There are a variety of OOD scenarios in the field of IR, so the OOD robustness of the model is likewise of wide interest. First, the query entered by the user may be unknown and of varying quality (Penha et al., 2022; Liu et al., 2023b). Then, search engine application scenario migration and incremental new documents will likewise bring in OOD data (Cai et al., 2023; Chen et al., 2023d; Thakur et al., 2021). Manual labeling of unseen data as well as retraining IR models incurs significant resource overhead (Nguyen et al., 2016; Thakur et al., 2021). Therefore, it is crucial how to efficiently train the IR model to have effective performance on unseen data.
In this context, studying the robustness of IR models on OOD data has received a lot of attention. OOD robustness measures the performance of an IR model on unseen queries and documents from different distributions of the training dataset. By introducing the test dataset $\tilde{\mathcal{D}}_{\mathrm{test}}$ in a new domain, we give a formal definition of OOD robustness in IR based on Definition 2.1.
Definition 2.0 (Out-of-distribution robustness of information retrieval).
Given an IR model $f_{\mathcal{D}_{\mathrm{train}}}$, an original dataset with training and test data, $\mathcal{D}_{\mathrm{train}}$ and $\mathcal{D}_{\mathrm{test}}$, drawn from the original distribution $\mathcal{G}$, along with a new test dataset $\tilde{\mathcal{D}}_{\mathrm{test}}$ drawn from the new distribution $\tilde{\mathcal{G}}$, and an acceptable error threshold $\delta$, for the top-$K$ ranking result, if
@@ 2.2.2. Out-of-distribution robustness
the model $f$ is considered $\delta$-robust against out-of-distribution data for metric $M$.
@@ 2.2.3. Adversarial robustness
In a competitive scenario, content providers may aim to promote their products or documents in rankings for specific queries (Kurland and Tennenholtz, 2022). This has provided a market for search engine optimization (SEO) and led to the development of attack techniques against search engines. Traditional attacks on search engines are generally called term spamming. They usually resort to stacking keywords to achieve a ranking boost. In recent years, with the development of deep learning, a number of neural approaches have emerged that attack through more imperceptible perturbations. As search engines evolve, defense methods against term spamming are maturing as well.
Adversarial robustness focuses on the stability of the IR model performance when imperceptible malicious perturbations are added to documents. By introducing a test dataset $\mathcal{D}^{\prime}_{\mathrm{test}}$ with adversarial examples, we give a formal definition of adversarial robustness in IR based on Definition 2.1.
@@
Given an IR model $f_{\mathcal{D}_{\mathrm{train}}}$ trained on training dataset $\mathcal{D}_{\mathrm{train}}$ with a corresponding testing dataset $\mathcal{D}_{\mathrm{test}}$, a new test dataset $\mathcal{D}^{\prime}_{\mathrm{test}}$ containing adversarial examples, and an acceptable error threshold $\delta$, for the top-$K$ ranking result, if
@@
where $\mathcal{D}_{\mathrm{test}}\cup D_{\mathrm{adv}}$ denotes injecting the set of all generated adversarial examples $D_{\mathrm{adv}}$ into the original test dataset, and then model $f$ is considered $\delta$-robust against adversarial examples for metric $M$.
In this paper, we pay special attention to two frequently studied types of robustness, i.e., adversarial robustness and OOD robustness. (i) In order to study adversarial robustness, existing work usually proceeds along two lines: adversarial attacks and adversarial defenses. We will discuss these lines in detail in Section 3. (ii) Depending on the types of OOD generalizability in IR models, the existing work can be categorized into the OOD generalizability on unseen documents and the OOD generalizability on unseen queries. We will discuss these directions in detail in Section 4. Research on performance variance mainly includes (Wu et al., 2022b; Zhang et al., 2013, 2014, 2022; Wang, 2009). Relatively little work has been done on this type of robustness.
@@ 3. Adversarial Robustness
The deep neural networks have been found vulnerable to adversarial examples that can produce misdirection with human-imperceptible perturbations (Goodfellow et al., 2015b; Ebrahimi et al., 2017). In the field of IR, deep learning-based models are also likely to inherit these adversarial vulnerabilities (Szegedy et al., 2014), which raises widespread concerns about the robustness of neural IR systems. Recall that we give the definition of adversarial robustness in IR elaborated in Section 2.2. Therefore, in this section, we present the specific work of adversarial attack and adversarial defense in adversarial robustness, respectively.
@@ 3.1. Overview
In IR, search engine optimization (SEO) has been around since the dawn of the world wide web (Gyöngyi and Garcia-Molina, 2005). This includes white-hat SEO (Goren et al., 2020), which modifies documents in good faith and within the rules and expectations of search engines to optimize the quality of web pages. In contrast, black-hat SEO (Castillo et al., 2011), maliciously exploiting loopholes of search engines, is used to get a site ranking higher in search results. Black-hat SEO creates a poor experience for the audience and is a common concern among site owners.
@@
SEO involves the creation of web pages with no real informational value, designed to manipulate search engine algorithms and distort search results. This deceptive practice lures web users to sites they would not typically visit, undermining the trust relationship between users and search engines and potentially damaging the search engines’ reputation (Baeza-Yates and Ribeiro-Neto, [n. d.]).
web spamming attack. When assessing textual relevance, search engines scrutinize the locations on a web page where query terms appear (Baeza-Yates and Ribeiro-Neto, [n. d.]). These locations are referred to as fields. Common text fields for a page include the document body, the title, the meta tags in the HTML header, and the page’s URL. Additionally, the anchor texts associated with URLs that point to the page are considered part of the page (anchor text field), as they often provide a succinct description of the page’s content. The terms in these text fields play a crucial role in determining the page’s relevance to a specific query, with different weights typically assigned to different fields. Term spamming refers to the manipulation of these text fields to make spam pages appear relevant for certain queries (Castillo et al., 2011).
web spamming detection. web spamming posed a serious threat to search engines in the early days. However, they are also easy to spot and suspect due to their simple implementation (Gyöngyi and Garcia-Molina, 2005). web spamming detection involves identifying and mitigating the practice of term spamming, where keywords are excessively used or manipulated on a web page to artificially inflate its relevance to specific search queries (Robertson and Walker, 1994). Various techniques have been developed to detect the term spamming.
@@
With the development of deep learning, neural networks are beginning to be widely used in IR models and have achieved excellent performance. Although traditional web spamming can also have a significant attack effect on neural IR models, this method is difficult to pose a threat due to its ease of detection. However, in the context of black-hat SEO, neural IR models are at risk of being attacked due to the inherent vulnerability inherited from neural networks. Therefore, adversarial attacks are beginning to be studied to expose vulnerability flaws in neural IR models in advance.
@@
The relationship between adversarial attacks and defenses is shown in Figure 5. Recently, adversarial robustness has begun to gain widespread attention. Below, we describe these efforts from several perspectives: benchmark datasets, adversarial attacks, and adversarial defense.
Table 1. Benchmark datasets for studying adversarial robustness. #Document denotes the number of documents in corpus; #Qtrain denotes the number of queries for training; #Qdev denotes the number of queries for development; #Qeval denotes the number of queries for evaluation.
Dataset #Document #Qtrain #Qdev #Qeval Reference
@@
Clueweb09-B (Clarke et al., 2009) 50M 150 - - (Wang et al., 2022a; Liu et al., 2024a)
@@
TREC DL19 (Craswell et al., 2020b) - - 43 - (Liu et al., 2022a; Wang et al., 2022a; Parry et al., 2024)
TREC DL20 (Craswell et al., 2021) - - 54 - (Parry et al., 2024)
TREC MB14 (Lin et al., 2013) - - 50 - (Liu et al., 2022a)
ASRC (Raifer et al., 2017) 1,279 - 31 - (Wu et al., 2022b; Goren et al., 2018)
Q-MS MARCO (Liu et al., 2023d) - - 4,000 - (Liu et al., 2023d)
Q-Clueweb09 (Liu et al., 2023d) - - 292 - (Liu et al., 2023d)
@@ 3.2. Benchmark Datasets
In this section, we present the datasets commonly used for studying adversarial robustness. Existing work on attacks and defenses against robustness has focused on experiments on existing IR datasets, as shown in Table 1. All datasets can be found in the BestIR benchmark.
@@
Some datasets in IR are often adapted for reuse by attack and defense methods as basic datasets. These include MS MARCO document/passage (Nguyen et al., 2016) and Clueweb-09B (Clarke et al., 2009). Some work (Wu et al., 2023a; Liu et al., 2023c, 2024a) performed experiments against attacks and defenses directly on these datasets. For example, prior work usually used the training set of the basic dataset to train an NRM as the target model (Wu et al., 2023a; Liu et al., 2024a). Meanwhile, the queries in the development set are used as target queries, and a portion of the documents in the ranked list of each query are sampled as target documents. Attacks on these target documents measure the performance of adversarial attack methods. We will discuss the specific evaluation methods in Section 3.3.
@@
Some competitions, such as TREC DL19 (Craswell et al., 2020b) and TREC DL20 (Craswell et al., 2021), have provided additional query collections for evaluation against the base dataset. Similarly, these query collections can be used to evaluate the effectiveness of attack methods. Queries in these datasets are often attacked as additional sets of targeted queries. For example, existing work usually trains NRM on the MS MARCO passage dataset and uses the 43 queries in TREC DL19 as the target queries to perform attacks (Chen et al., 2023c; Liu et al., 2022a).
@@
Some research has adapted the above basic datasets to construct new datasets that can be used directly to perform attacks or evaluate defenses. For example, ASRC (Raifer et al., 2017) is based on documents in Clueweb, which are manually modified to generate new adversarial samples for evaluating the model’s adversarial defense abilities. Existing work also used it to study the effects of manual manipulation of documents on the environment IR systems (Vasilisky et al., 2023; Goren et al., 2018). To perform the topic-oriented attack, Liu et al. (2023d) constructed query groups on the same topic based on ORCAS (Craswell et al., 2020a) and the TREC 2012 web Track (Clarke et al., 2009) as a complement to MS MARCO document and Clueweb-09b, respectively. DARA (Chen et al., 2023b) is a dataset for detecting adversarial ranking attacks and includes two types of detection tasks for adversarial documents.
3.3. Adversarial Attack
@@ 3.3.1. What are the differences between IR attacks and CV/NLP attacks?
Adversarial attacks have undergone significant development in the fields of NLP and CV (Ebrahimi et al., 2017; Goodfellow et al., 2015b), where the most emblematic forms of attack are typically directed at image retrieval and text classification tasks. However, the landscape of adversarial attacks assumes a different contour in the realm of IR: (i) Compared with image retrieval attacks, the IR attacks need to maintain semantic consistency of the perturbed document with the original document by considering the textual semantic similarity, rather than pixel-level perturbations within a fixed range in continuous space; and (ii) Inspired by black-hat SEO, the goal of IR attacks is to inject imperceptible perturbations within a document to improve its ranking for one or multiple specific queries within the entire candidate set or corpus, not inducing classification errors by the model.
@@
where $\pi_{f}\left(q,d\oplus p\right)$ denotes the ranking position of the perturbed document $d\oplus p$ in the ranked list generated by $f$ with respect to query $q$. The $\operatorname{Sim}\left(\cdot\right)$ function measures the similarity between the adversarial example and the original document, both textually as well as semantically. $\lambda$ is a regularization parameter used to balance two goals: keeping the adversarial samples as close as possible to the original document, while allowing the adversarial samples to be ranked as high as possible. Ideally, the adversarial sample $d\oplus p$ should preserve the original semantics of $d$, and be imperceptible to human judges yet misleading to the neural IR models.

### v2

@@ 2.2. Taxonomy of Robustness in IR
In IR, robustness threats come in different flavors, including IID robustness (Wu et al., 2022b), OOD robustness (Thakur et al., 2021), and adversarial robustness (Wu et al., 2022b). A high-level taxonomy of robustness in IR is shown in Figure 4.
2.2.1. IID robustness
Typically, the performance of IR models is first represented by their overall performance on IID data. Recently, it has been recognized that performance stability across IID queries may be compromised when we try to improve the average retrieval effectiveness across all queries (Zhang et al., 2013). Therefore, a robust neural IR model should not only have good retrieval performance on the overall testing queries, but also ensure that the performance on individual queries is not too bad. We give a formal definition of IID robustness in IR based on Definition 2.1.
Definition 2.0 (IID robustness of information retrieval).
Let the following be given: an IR model $f_{\mathcal{D}_{\mathrm{train}}}$ trained on training dataset $\mathcal{D}_{\mathrm{train}}$ with a corresponding IID test dataset $\mathcal{D}_{\mathrm{test}}$, and an acceptable error threshold $\delta$, for the top-$K$ ranking result. If
@@ 2.2.1. IID robustness
where $\mathrm{Var}(\cdot)$ is the variance of the ranking performance of the IR model $f_{\mathcal{D}_{\mathrm{train}}}$ on $\mathcal{D}_{\mathrm{test}}$, then the model $f$ is considered to be $\delta$-robust in terms of IID data for metric $M$.
2.2.2. OOD robustness
IR models need to cope with a constant stream of unseen data (Wu et al., 2022b). The key behind addressing this challenge is how to adapt the model to new data out of the familiar distribution. There are a variety of OOD scenarios in IR, so the OOD robustness of the model is of broad interest. First, the query entered by the user may be unknown and of varying quality (Penha et al., 2022; Liu et al., 2023b). Then, search engine application scenario migration and incremental new documents will likely bring in OOD data (Cai et al., 2023; Chen et al., 2023d; Thakur et al., 2021). Manual labeling of unseen data as well as retraining IR models incurs significant resource overhead (Nguyen et al., 2016; Thakur et al., 2021). Therefore, it is crucial how to efficiently train IR models to achieve effective performance on unseen data.
OOD robustness measures the performance of an IR model on unseen queries and documents from distributions that differ from the training dataset. By introducing the test dataset $\tilde{\mathcal{D}}_{\mathrm{test}}$ in a new domain, we give a formal definition of OOD robustness in IR based on Definition 2.1.
Definition 2.0 (OOD robustness of information retrieval).
Let the following be given: an IR model $f_{\mathcal{D}_{\mathrm{train}}}$, an original dataset with training and test data, $\mathcal{D}_{\mathrm{train}}$ and $\mathcal{D}_{\mathrm{test}}$, drawn from the original distribution $\mathcal{G}$, along with a new test dataset $\tilde{\mathcal{D}}_{\mathrm{test}}$ drawn from the new distribution $\tilde{\mathcal{G}}$, and an acceptable error threshold $\delta$, for the top-$K$ ranking result. If
@@ 2.2.2. OOD robustness
the model $f$ is considered to be $\delta$-robust against OOD data for metric $M$.
@@ 2.2.3. Adversarial robustness
In a competitive scenario, content providers may aim to promote their products or documents in rankings for specific queries (Kurland and Tennenholtz, 2022). This has provided a market for SEO and has led to the development of attack techniques against search engines. Traditional attacks against search engines are generally called term spamming. They usually resort to stacking keywords to achieve a ranking boost. In recent years, with the development of deep learning, a number of neural approaches have emerged that attack through more imperceptible perturbations. As search engines evolve, defense methods against term spamming are maturing as well. Adversarial robustness focuses on the stability of the IR model performance when imperceptible malicious perturbations are added to documents. By introducing a test dataset $\mathcal{D}^{\prime}_{\mathrm{test}}$ with adversarial examples, we give a formal definition of adversarial robustness in IR based on Definition 2.1.
@@
Let the following be given: an IR model $f_{\mathcal{D}_{\mathrm{train}}}$ trained on training dataset $\mathcal{D}_{\mathrm{train}}$ with a corresponding testing dataset $\mathcal{D}_{\mathrm{test}}$, a new document set $D_{\mathrm{adv}}$ containing adversarial examples, and an acceptable error threshold $\delta$, for the top-$K$ ranking result. If
@@
where $\mathcal{D}_{\mathrm{test}}\cup D_{\mathrm{adv}}$ denotes injecting the set of all generated adversarial examples $D_{\mathrm{adv}}$ into the original test dataset, then model $f$ is considered to be $\delta$-robust against adversarial examples for metric $M$.
Relatively little work has been done on IID robustness. Research on this type of robustness mainly includes (Wu et al., 2022b; Zhang et al., 2013, 2014, 2022; Wang, 2009). In this paper, we pay special attention to the two other notions of robustness depicted in Figure 4, i.e., adversarial robustness and OOD robustness. Existing work on adversarial robustness usually proceeds along two lines: adversarial attacks and adversarial defenses. We will discuss these lines in detail in Section 3. Depending on the types of OOD generalizability in IR models, the existing work can be categorized into the OOD generalizability on unseen documents and the OOD generalizability on unseen queries. We will discuss these directions in detail in Section 4.
@@ 3. Adversarial Robustness
Deep neural networks have been found to be vulnerable to adversarial examples that can produce misdirection with human-imperceptible perturbations (Goodfellow et al., 2015b; Ebrahimi et al., 2017). In IR, deep learning-based models are also likely to inherit these adversarial vulnerabilities (Szegedy et al., 2014), which raises concerns about the robustness of neural IR systems. Building on the definitions of adversarial robustness in IR provided in Section 2.2, we survey work on adversarial robustness, focusing specifically on adversarial attacks and adversarial defenses.
@@ 3.1. Overview
In IR, search engine optimization (SEO) has been around since the dawn of the world wide web (Gyöngyi and Garcia-Molina, 2005). This includes white-hat SEO (Goren et al., 2020), which modifies documents in good faith and within the rules and expectations of search engines to optimize the quality of web pages. In contrast, black-hat SEO (Castillo and Davison, 2011), maliciously exploiting loopholes of search engines, is used to get a site ranking higher in search results.
@@
SEO involves the creation of web pages with no real informational value, designed to manipulate search engine algorithms and distort search results. This deceptive practice lures web users to sites they would not typically visit, undermining the trust relationship between users and search engines and potentially damaging the search engine’s reputation (Baeza-Yates and Ribeiro-Neto, 1999).
Web spamming attack.
When assessing textual relevance, search engines scrutinize the locations on a web page where query terms appear (Baeza-Yates and Ribeiro-Neto, 1999). These locations are referred to as fields. Common text fields for a page include the document body, the title, the meta tags in the HTML header, and the page’s URL. Additionally, the anchor texts associated with URLs that point to the page are considered part of the page (anchor text field), as they often provide a succinct description of the page’s content. The terms in these text fields play a crucial role in determining the page’s relevance to a specific query, with different weights typically assigned to different fields. Term spamming refers to the manipulation of these text fields to make spam pages appear relevant for certain queries (Castillo and Davison, 2011).
Web spamming detection.
Web spamming posed a serious threat to search engines in the early days. However, they are also easy to spot and suspect due to their simple implementation (Gyöngyi and Garcia-Molina, 2005). Web spamming detection involves identifying and mitigating the practice of term spamming, where keywords are excessively used or manipulated on a web page to artificially inflate its relevance to specific search queries (Robertson and Walker, 1994). Various techniques have been developed to detect term spamming.
@@
With the development of deep learning, neural networks are beginning to be widely used in IR models and have achieved excellent performance. Although traditional web spamming can also have a significant attack effect on neural IR models, this method does not really pose a threat due to its ease of detection. However, in the context of black-hat SEO, neural IR models are at risk of being attacked due to the inherent vulnerability inherited from neural networks. Therefore, adversarial attacks are beginning to be studied to expose vulnerability flaws in neural IR models in advance.
AttackDefenseTop-$K$ resultschangeidentify flawsstabilizegstrengthen models
@@
The relationship between adversarial attacks and defenses is shown in Figure 5. Recently, adversarial robustness has begun to gain widespread attention. Below, we describe these efforts from several perspectives: benchmark datasets, adversarial attacks, and adversarial defenses.
Table 1. Benchmark datasets for studying adversarial robustness. #Documents denotes the number of documents in the corpus; #Qtrain denotes the number of queries available for training; #Qdev denotes the number of queries available for development; #Qeval denotes the number of queries available for evaluation.
Dataset #Documents #Qtrain #Qdev #Qeval References
@@
Clueweb09-B (Clarke et al., 2009) 50M 150 (Wang et al., 2022a; Liu et al., 2024a)
@@
TREC DL19 (Craswell et al., 2020b) 43 (Liu et al., 2022a; Wang et al., 2022a; Parry et al., 2024)
TREC DL20 (Craswell et al., 2021) 54 (Parry et al., 2024)
TREC MB14 (Lin et al., 2013) 50 (Liu et al., 2022a)
ASRC (Raifer et al., 2017) 1,279 31 (Wu et al., 2022b; Goren et al., 2018)
Q-MS MARCO (Liu et al., 2023d) 4,000 (Liu et al., 2023d)
Q-Clueweb09 (Liu et al., 2023d) 292 (Liu et al., 2023d)
@@ 3.2. Benchmark Datasets
In this section, we present the datasets commonly used for studying adversarial robustness. Prior work on attacks and defenses against robustness has focused on experiments on previusly published IR datasets, as shown in Table 1. All datasets can be found in the BestIR benchmark; see Appendix B for an overview of the resources collected in BestIR.
@@
Some datasets in IR are adapted for reuse by attack and defense methods as basic datasets. These include MS MARCO document/passage (Nguyen et al., 2016) and Clueweb-09B (Clarke et al., 2009). Some work (Wu et al., 2023a; Liu et al., 2023c, 2024a) performed experiments against attacks and defenses directly on these datasets. For example, prior work usually used the training set of the basic dataset to train an NRM as the target model (Wu et al., 2023a; Liu et al., 2024a); the queries in the development set are used as target queries, and a portion of the documents in the ranked list of each query is sampled as target documents. Attacks on these target documents measure the performance of adversarial attack methods. We will discuss the specific evaluation methods in Section 3.3.
@@
Some collaborative benchmarks, such as TREC DL19 (Craswell et al., 2020b) and TREC DL20 (Craswell et al., 2021), have provided additional query collections for evaluation against the base dataset. Similarly, these query collections can be used to evaluate the effectiveness of attack methods. Queries in these datasets are often attacked as additional sets of targeted queries. For example, existing work usually trains an NRM on the MS MARCO passage dataset and uses the 43 queries in TREC DL19 as the target queries to perform attacks (Chen et al., 2023c; Liu et al., 2022a).
@@
Some research has adapted the above datasets to construct new datasets that can be used directly to perform attacks or evaluate defenses. For example, ASRC (Raifer et al., 2017) is based on documents in Clueweb, which are manually modified to generate new adversarial samples for evaluating the model’s adversarial defense abilities. Existing work also used it to study the effects of manual manipulation of documents on IR systems (Vasilisky et al., 2023; Goren et al., 2018). To perform a topic-oriented attack, Liu et al. (2023d) constructed query groups on the same topic based on ORCAS (Craswell et al., 2020a) and the TREC 2012 Web Track (Clarke et al., 2009) as a complement to MS MARCO document and Clueweb-09b, respectively. DARA (Chen et al., 2023b) is a dataset for detecting adversarial ranking attacks and includes two types of detection task for adversarial documents.
3.3. Adversarial Attacks
@@ 3.3.1. What are the differences between IR attacks and CV/NLP attacks?
Adversarial attacks have undergone significant development in the fields of CV and NLP (Ebrahimi et al., 2017; Goodfellow et al., 2015b), where attacks are typically directed at image retrieval and text classification tasks, respectively. However, the landscape of adversarial attacks differs in IR: (i) Compared with image retrieval attacks, the IR attacks need to maintain semantic consistency of the perturbed document with the original document by considering the textual semantic similarity, rather than pixel-level perturbations within a fixed range in continuous space; and (ii) Inspired by black-hat SEO, the goal of IR attacks is to inject imperceptible perturbations within a document to improve its ranking for one or multiple specific queries within the entire candidate set or corpus, not inducing classification errors by the model.
@@
where $\pi_{f}\left(q,d\oplus p\right)$ denotes the ranking position of the perturbed document $d\oplus p$ in the ranked list generated by $f$ with respect to query $q$. The $\operatorname{Sim}\left(\cdot\right)$ function measures the similarity between the adversarial example and the original document, both textually as well as semantically. $\lambda$ is a regularization parameter used to balance two goals: keeping the adversarial samples as close as possible to the original document, while allowing adversarial samples to be ranked as high as possible. Ideally, the adversarial sample $d\oplus p$ should preserve the original semantics of $d$, and be imperceptible to human judges yet misleading to the neural IR models.